OpenStudy (anonymous):

Suppose y varies directly with x. If x is 30 when y is 10, what is x when y is 9.

OpenStudy (espex):

There is a direct relation between x and y, so you know that whatever y does, x does proportionally. \[\frac{x}{y} \rightarrow \frac{30}{10}=\frac{x}{9}\] Simply solve for x.

OpenStudy (anonymous):

How do i solve for x?